### v2

# David Marín111david.marin”AT”uab.cat

Abstract
We study degenerate singular points of planar vector fields inside a (degenerate) flow box. These kind of singularities are called fake saddles and their linear parts are always zero. We characterize fake saddles with non-zero second-order jet and we give the first term of a uniform asymptotic expansion of the Poincaré map between two transverse sections to their corresponding singular fiber, determining its stability.
Following [6, 7] we define a fake saddle as a singular point having exactly two separatrices and two hyperbolic sectors, see Figure 1. Another way to think about a fake saddle is that the corresponding vector field can be put in a form that is like a degenerate flow box, i.e. near the singularity the phase portrait consists of parallel fibers, all but one of which have no singular points, and the singular fiber (the union of the singular point with its two separatrices) has a semi-stable equilibrium point, see [2]. The local phase portrait of a singularity with non-zero linear part (for the nilpotent case see for instance [1, Theorem 3.5]) is well known and prevents the singular point from being a fake saddle. The first non-trivial (and generic) case of false saddle occurs when the second-order jet is non-zero. Our goal is to characterize such generic fake saddles and also to study their transition maps between two transverse sections to the singular fiber in order to know its attractive or repulsive character. The techniques we use allow to give the first term of an asymptotic expansion which is uniform with respect to any parameters of the fake saddle.
Figure 1: Fake saddle and the associated transition maps.
There exist nilpotent singular points that are not fake saddles for which $y=0$ is a separatrix having a transition map defined only on one side $\pm y>0$. For instance, the vector field
$(\lambda y+x^{2}+y^{2})\partial_{x}-\frac{xy}{\lambda}\partial_{y},\quad\lambda>0,$
with Darboux first integral $|y|^{2\lambda}\left(x^{2}+\frac{\lambda}{\lambda+1}y^{2}+\frac{2\lambda^{2}}{1+2\lambda}y\right)$, has a well-defined transition map along $y=0$ only for $y>0$. The singularity at the origin has four characteristic orbits (or separatrices) $\{\pm x>0,y=0\}$, $\{\pm x>0,x^{2}+\frac{\lambda}{\lambda+1}y^{2}+\frac{2\lambda^{2}}{1+2\lambda}y=0\}$, separating four hyperbolic sectors. $\square$
This work is mainly motivated by the paper [2] which provides a normal form for generic fake saddles that we recall in the sequel. First we can locate the fake saddle at the origin $(0,0)$. Secondly, if the vector field is at least of class $\mathscr{C}^{3}$ there exists a local diffeomorphism that rectifies the germ of singular fiber, so that it is contained in the line $y=0$ in a certain coordinate system $(x,y)$, see [2, Lemma 2.1]. Thanks to the hypothesis that the second-order jet of the fake saddle is non-zero, after a suitable rescaling of the coordinates we can assume that the fake saddle is defined by a system of differential equations
$\left\{\begin{array}[]{l}\dot{x}=x^{2}+axy+y^{2}+\mathrm{O}_{3}(x,y),\\ \dot{y}=y\big(cx+by+\mathrm{O}_{2}(x,y)\big),\end{array}\right$.
where $a,b,c\in\mathbb{R}$. $f_{1}(0,0)=f_{2}(0,0)=1$. This normal form motivates us to consider a smooth family of planar vector fields $\{X_{\mu}\}_{\mu\in\Omega\subset\mathbb{R}^{n}}$ having the form
$X_{\mu}(x,y)=(x^{2}f_{1}(x,y;\mu)+a(\mu)xy+y^{2}f_{2}(x,y;\mu))\partial_{x}+\big(xg_{1}(x,y;\mu)+yg_{2}(y;\mu)\big)y\partial_{y},$ (1)
where $f_{1}(x,y;\mu)$, $f_{2}(x,y;\mu)$, $g_{1}(x,y;\mu)$ and $g_{2}(y;\mu)$ are $\mathscr{C}^{K}$-functions fulfilling $f_{1}(0,0;\mu)=f_{2}(0,0;\mu)=1$. We define the following invariants
$a(\mu),\quad b(\mu)=g_{2}(0;\mu),\quad c(\mu)=g_{1}(0,0;\mu)$ (2)
of the family (1) and consider the associated value $d(\mu):=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}$ which will play a key role in the sequel.
In this paper we extend the main theorems of [2] and we give simpler proofs of them using the results stated in [4, 5]. Our first result gives a characterization of generic fake saddles inside the family (1), generalizing [2, Theorem A] which only treats the case $a=0$.
Theorem 2.
If the invariants $(a,b,c)$ given in (2) of a vector field $X$ in the family (1) do not belong to $\{d=0\}\cap\{a^{2}-b^{2}=4\}$ then the origin is a fake saddle if and only if, either $d>0$, or $c=1$ and $a=b$. In both situations after blowing up the origin we have a single singular point on the exceptional divisor, which is a hyperbolic saddle of hyperbolicity ratio 1c1-c in the first one and a semi-hyperbolic saddle in the second one.
Proof.
We have that
$Y=\frac{(u,uv)^{*}X}{u}=P(u,v)u\partial_{u}+Q(u,v)v\partial_{v}$
with $Q(0,v)=-v^{2}+(b-a)v+c-1$. The point $(u,v)=(0,0)$ is always a singular point of $Y$. If $d<0$ then $Y$ has two other singular points on the exceptional divisor $u=0$ with at least one non-zero eigenvalue. If $(a,b,c)\in\{d=0\}\setminus(\{c=1,a=b\}\cup\{a^{2}-b^{2}=4\})$ then $Y$ has another double singular point at (u,v)=(0,(ba)/2)(u,v)=(0,(b-a)/2) with a non-zero eigenvalue. In these situations the transition map $\Pi_{\alpha}^{\omega}:\{x=\alpha\}\to\{x=\omega\}$ is not well defined so that the origin is not a fake saddle of $X$. The remaining assertions are easy to check. ∎
The hypothesis $(a,b,c)\notin\{d=0\}\cap\{a^{2}-b^{2}=4\}$ in Theorem 2 (which is always verified when $a=0$) can not be removed as the following example shows.
For each integer $n\geq 3$ we consider the vector field
$X_{n}=(x+y)^{2}\partial_{x}+y^{n}\partial_{y}$
having a degenerate singularity at the origin with invariants $(a,b,c)=(2,0,0)\in\{d=0\}\cap\{a^{2}-b^{2}=4\}$.
It can be checked that in the resolution of singularities of $X_{3}$ appears a saddle-node whose weak separatrix is not the strict transform of $y=0$ and meets transversely the exceptional divisor. Consequently the origin is not a fake saddle for the vector field $X_{3}$.
On the other hand, the blowup of the vector field $X_{4}$
$Y_{0}=\frac{(v,uv)^{*}X_{4}}{v}=(-(u+1)^{2}+u^{3}v^{2})u\partial_{u}+(u+1)^{2}v\partial_{v}$
has two singular points on the exceptional divisor $v=0$: $(u,v)=(0,0)$ which is a hyperbolic saddle and (u,v)=(1,0)(u,v)=(-1,0) which is degenerated. Moreover,
$Y_{1}=(u-1,v)^{*}Y_{0}=(u^{2}+v^{2}-u^{3}-4uv^{2}+6u^{2}v^{2}-4u^{3}v^{2}+u^{4}v^{2})\partial_{u}+u^{2}v\partial_{v}$
has invariants $(a,b,c)=(0,0,0)$, so that $d=4$ and the transition map along $v=0$ is well-defined for $Y_{1}$. Thus, the origin is a fake saddle of $X_{4}$, see Figure 2. This example (for which $a_{1,1}:=a=2$, $h_{0,1}:=b=0$ and $h_{1,0}:=c=0$) contradicts the following claim of [2]: “it is readily seen that a necessary condition for the origin to be a fake singularity on the singular fiber $y=0$ is that […] either $(h_{0,1}-a_{1,1})^{2}+4(h_{1,0}-1)<0$, or that $h_{1,0}=1$ and $h_{0,1}=a_{1,1}$.”
Figure 2: Numerical solutions of $X_{4}=(x+y)^{2}\partial_{x}+y^{4}\partial_{y}$.
Moreover, since $a=b=0$, according to Theorem 4 below, the derivative at $v=0$ of the transition map $\Pi_{\alpha}^{\omega}(v)$ of $Y_{1}$ is
$\exp\mathrm{PV}\int_{\alpha}^{\omega}\frac{u^{2}}{u^{2}-u^{3}}du=\left|\frac{1-\alpha}{1-\omega}\right|,$
so that it is contractive or repulsive on both sides $\pm v>0$. Notice that the transition map $\Pi_{-1}^{+1}(y)$ of $X_{4}$ is contractive on one side $y>0$ and repulsive on the other side $y<0$, see Figure 2. $\square$
Let $\Omega$ be an open subset of $\mathbb{R}^{n}$ and let $W$ any subset of $\Omega$. Recall (cf. [4, Definition 1.2]) that a $\mathscr{C}^{K}$-function $f(s;\mu)$ defined in the intersection of $(0,+\infty)\times\Omega$ with an open neighborhood of $\{0\}\times\Omega\subset\mathbb{R}^{n+1}$ belongs to the flat class $\mathcal{F}_{L}^{K}(W)$ if for each $\mu_{0}\in W$ and $\nu=(\nu_{0},\ldots,\nu_{n})\in\mathbb{N}^{n+1}$ with $|\nu|=\nu_{0}+\cdots+\nu_{n}\leq K$ there exists $\varepsilon>0$ such that $|\partial_{s}^{\nu_{0}}\partial_{\mu_{1}}^{\nu_{1}}\cdots\partial_{\mu_{n}}^{\nu_{n}}f(s;\mu)|\leq Cs^{L-\nu_{0}}$ for every $s\in(0,\varepsilon)$ and $\mu\in\mathbb{B}_{\varepsilon}(\mu_{0})\cap\Omega$.
We consider the transverse sections $\Sigma_{\star}=\{x=\star\}$ to $y=0$, for $\star\in\{\alpha,\omega\}$, where $\alpha<0<\omega$. We assume that $f_{1}(x,0)>0$ for all $x\in[\alpha,\omega]$ and we consider the transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$. Our second result extends [2, Theorem B], which only treats the case $a=b=0$ when the transition map is restricted to $y>0$.
Theorem 4.
Assume that $\mu_{0}\in\Omega$ and the invariants (2) of the vector field $X_{\mu_{0}}$ in (1) belong to $\{d=4(1-c)-(a-b)^{2}>0\}$. If $\mu\approx\mu_{0}$ then the origin is a fake saddle of $X_{\mu}$ and its transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$ satisfies $\Pi_{\alpha}^{\omega}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{K}(\{d>0\})$ on $\pm y\geq 0$, where $\epsilon>0$ and
$\gamma_{\pm}(\mu)=\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0;\mu)}{xf_{1}(x,0;\mu)}dx\pm\frac{\pi(2b(\mu)-c(\mu)(a(\mu)+b(\mu)))}{\sqrt{d(\mu)}}.$
Clearly the sign of $\gamma_{\pm}(\mu)$ determines the stability of the singular fiber $y=0$ on the side $\pm y>0$. In the case $a=b=0$ the two values $\gamma_{\pm}(\mu)$ coincide with the principal value of the integral. Otherwise the stability of the two sides $\pm y>0$ can be different, see Example Fake saddles and their transition maps 00footnotetext: 2020 AMS Subject Classification: 34D10, 34D20; Secondary: 34C05. 00footnotetext: Key words and phrases: singularities, Poincaré and Dulac transition maps, asymptotic expansion, uniform flatness, stability 00footnotetext: This work has been partially funded by the Ministry of Science, Innovation and Universities of Spain through the grant PID2021-125625NB-I00 and by the Agency for Management of University and Research Grants of Catalonia through the grant 2021SGR01015. . $\square$
It is worth to be noticed that the writing of $\Pi_{\alpha}^{\omega}(y;\mu)$ in the statement implies that it is of the form $e^{\gamma_{\pm}(\mu)}y+\mathrm{o}(y)$ but also that the remainder term $\mathrm{o}(y)$ is uniform with respect to the parameter $\mu$. This uniformity allows to address cyclicity problems and not just study stability. For clarity in the exposition we will omit the dependence on $\mu$ when it is not essential.
Using that $f_{1}(0,0)=1$ and $g_{1}(0,0)=c$ we can compute the previous principal value as
$\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx:=\lim\limits_{\varepsilon\to 0^{+}}\int_{(\alpha,\omega)\setminus(-\varepsilon,\varepsilon)}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx=c\log\left|\frac{\omega}{\alpha}\right|+\int_{\alpha}^{\omega}\left(\frac{g_{1}(x,0)}{f_{1}(x,0)}-c\right)\frac{dx}{x}.$
$\square$
Proof.
Let us compute first the value $\gamma_{+}$, i.e. we assume that $y\geq 0$. Consider the charts $(x,y)=\pi_{\pm}(u,v)=(\pm u(1-v),uv)$ of the blow-up of the origin, see Figure 3.
$\Sigma_{\alpha}$$\Sigma_{0}$$\Sigma_{\omega}$$\{x=y\}$$\{x=-y\}$xxyy
Figure 3: Blowup of the fake saddle at the origin.
Notice that $v=\frac{y}{y\pm x}$ is a coordinate on the exceptional divisor $u=0$. We assume that $u,v\geq 0$, we write
$\frac{\pi_{\pm}^{*}X}{u}=P_{\pm}(u,v)u\partial_{u}+Q_{\pm}(u,v)v\partial_{v}$
and define
$X_{\pm}(x,y)=P_{1}^{\pm}(x,y)x\partial_{x}+P_{2}^{\pm}(x,y)y\partial_{y}$
with
$P_{1}^{-}(x,y)=Q_{-}(y,x),\quad P_{2}^{-}(x,y)=P_{-}(y,x),\quad P_{1}^{+}(x,y)=P_{+}(x,y),\quad P_{2}^{+}(x,y)=Q_{+}(x,y).$
It can be checked that $\lambda_{\pm}=-\frac{P_{2}^{\pm}(0,0)}{P_{1}^{\pm}(0,0)}>0$, so that $(0,0)$ is a hyperbolic saddle of $X_{\pm}$. Consider the Dulac map $D_{\pm}$ of $X_{\pm}$ between the parametrized transverse sections
$\sigma_{1}^{-}(s)=\Big(\frac{s}{-\alpha+s},-\alpha+s\Big),\quad\sigma_{2}^{-}(s)=(1,s),\quad\sigma_{1}^{+}(s)=(s,1),\quad\sigma_{2}^{+}(s)=\Big(\omega+s,\frac{s}{\omega+s}\Big).$ (3)
We know from [4, Theorem A] that $D_{\pm}(s;\mu)=s^{\lambda_{\pm}(\mu)}(\Delta_{00}^{\pm}(\mu)+\mathcal{F}_{\epsilon})$ for some $\epsilon>0$, where we write $\mathcal{F}_{\epsilon}$ instead of $\mathcal{F}_{\epsilon}^{K}(\{\lambda_{\pm}>0\})$ for brevity. Then, using [4, Lemma A.2], we can write
$\Pi_{\alpha}^{\omega}(y)=(D_{+}\circ D_{-})(y)=\Big(y^{\lambda_{-}}(\Delta_{00}^{-}+\mathcal{F}_{\epsilon})\Big)^{\lambda_{+}}(\Delta_{00}^{+}+\mathcal{F}_{\epsilon})=\Delta_{00}y+\mathcal{F}_{1+\epsilon},$
with $\Delta_{00}=(\Delta_{00}^{-})^{\lambda_{+}}\Delta_{00}^{+}$. Following [5], to compute the coefficients $\Delta_{00}^{\pm}$ we must introduce some auxiliary functions:
$R_{12}^{-}(y)$ $=\frac{P_{1}^{-}}{P_{2}^{-}}(0,y)=\frac{g(-y,0)}{f(-y,0)}-1,$
$R_{21}^{-}(x)$ $=\frac{P_{2}^{-}}{P_{1}^{-}}(x,0)=\frac{1}{1-c}\left[\frac{\left(-a+b+c-2\right)x^{2}+\left(a-c+2\right)x-1}{\left(a-b-c+2\right)x^{2}+\left(-a+b+2c-2\right)x+1-c}\right],$
$R_{12}^{+}(y)$ $=\frac{P_{1}^{+}}{P_{2}^{+}}(0,y)=\frac{1}{1-c}\left[\frac{\left(-a+b-c+2\right)y^{2}+\left(a+c-2\right)y+1}{\left(a-b+c-2\right)y^{2}+\left(-a+b-2c+2\right)y+c-1}\right],$
$R_{21}^{+}(x)$ $=\frac{P_{2}^{+}}{P_{1}^{+}}(x,0)=\frac{g(x,0)}{f(x,0)}-1.$
Notice that $\lambda:=\lambda_{+}=1-c>0$ and and $\lambda_{-}=\frac{1}{\lambda_{+}}$. We have that
$\log L_{1}^{-}(u)$ $:=\int_{0}^{u}\left(R_{12}^{-}(y)+\frac{1}{\lambda_{-}}\right)\frac{dy}{y}=\int_{0}^{-u}\left(\frac{g(x,0)}{f(x,0)}-c\right)\frac{dx}{x},$
$\log L_{2}^{-}(u)$ $:=\int_{0}^{u}\left(R_{21}^{-}(x)+{\lambda_{-}}\right)\frac{dx}{x}=\frac{1}{1-c}\int_{0}^{u}\frac{\Big(c\left(a-b-c+2\right)x-\left(ac-c^{2}-b+c\right)\Big)dx}{\left(a-b-c+2\right)x^{2}-\left(a-b-2c+2\right)x-\left(c-1\right)},$
$\log L_{1}^{+}(u)$ $:=\int_{0}^{u}\left(R_{12}^{+}(y)+\frac{1}{\lambda_{+}}\right)\frac{dy}{y}=\frac{1}{1-c}\int_{0}^{u}\frac{\Big(c\left(a-b+c-2\right)y-\left(ac+c^{2}-b-c\right)\Big)dy}{\left(a-b+c-2\right)y^{2}-\left(a-b+2c-2\right)y+\left(c-1\right)},$
$\log L_{2}^{+}(u)$ $:=\int_{0}^{u}\left(R_{21}^{+}(x)+{\lambda_{+}}\right)\frac{dx}{x}=\int_{0}^{u}\left(\frac{g(x,0)}{f(x,0)}-c\right)\frac{dx}{x}.$
From (3) we compute the partial derivatives $\sigma_{ijk}^{\pm}=\partial_{s}^{k}\sigma_{ij}^{\pm}(0)$ of the parametrizations $\sigma_{i}^{\pm}(s)=(\sigma_{i1}^{\pm}(s),\sigma_{i2}^{\pm}(s))$:
$\sigma^{-}_{111}=-1/\alpha,\quad\sigma^{-}_{120}=-\alpha,\quad\sigma^{-}_{210}=1,\quad\sigma^{-}_{221}=1,\quad\sigma_{111}^{+}=1,\quad\sigma_{120}^{+}=1,\quad\sigma_{210}^{+}=\omega,\quad\sigma_{221}^{+}=1/\omega.$
Then, according to [5, Theorem A] and Remark Fake saddles and their transition maps 00footnotetext: 2020 AMS Subject Classification: 34D10, 34D20; Secondary: 34C05. 00footnotetext: Key words and phrases: singularities, Poincaré and Dulac transition maps, asymptotic expansion, uniform flatness, stability 00footnotetext: This work has been partially funded by the Ministry of Science, Innovation and Universities of Spain through the grant PID2021-125625NB-I00 and by the Agency for Management of University and Research Grants of Catalonia through the grant 2021SGR01015. , we obtain the following expression:
$\Delta_{00}$ $=(\Delta_{00}^{-})^{\lambda_{+}}\Delta_{00}^{+}=\frac{\sigma_{111}^{-}(\sigma_{120}^{-})^{\lambda}(L_{2}^{-}(\sigma_{210}^{-}))^{\lambda}(\sigma_{111}^{+})^{\lambda}\sigma_{120}^{+}L_{2}^{+}(\sigma_{210}^{+})}{L_{1}^{-}(\sigma_{120}^{-})(\sigma_{221}^{-})^{\lambda}\sigma_{210}^{-}(L_{1}^{+}(\sigma_{120}^{+}))^{\lambda}\sigma_{221}^{+}(\sigma_{210}^{+})^{\lambda}}=\frac{(-\alpha)^{-1+\lambda}L_{2}^{+}(\omega)}{\omega^{-1+\lambda}L_{1}^{-}(-\alpha)}\left(\frac{L_{2}^{-}(1)}{L_{1}^{+}(1)}\right)^{\lambda}$
$=\exp\left(\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx+\gamma_{0}\right),$
where $\gamma_{0}:=\lambda\log\left(\frac{L_{2}^{-}(1)}{L_{1}^{+}(1)}\right)$. Let us prove now that
$\gamma_{0}=-\frac{\pi(c(a+b)-2b)}{\sqrt{d}}.$
With this aim notice that
$\log L_{1}^{+}(u;a,b,c)=\log L_{2}^{-}(u;-a,-b,c)=\alpha(u;a,b,c)+\beta(u;a,b,c)$
where
$\alpha(u;a,b,c)=\frac{c}{2(1-c)}\log\left(\frac{1-c+(a-b+2c-2)u+(-a+b-c+2)u^{2}}{1-c}\right)$
and
$\beta(u;a,b,c)={\scriptstyle-\frac{((a+b)c-2b)}{(1-c)\sqrt{d}}}\left[\arctan\left(\frac{2(1-c)+b-a+2(a-b+c-2)u}{\sqrt{d}}\right)-\arctan\left(\frac{2(1-c)+b-a}{\sqrt{d}}\right)\right].$
Since $\alpha(1;a,b,c)=-\frac{c}{2(1-c)}\log(1-c)$ we obtain that
$\log\left(\frac{L_{2}^{-}(1)}{L_{1}^{+}(1)}\right)=\beta(1;-a,-b,c)-\beta(1;a,b,c)=\frac{((a+b)c-2b)}{(1-c)\sqrt{d}}F(a,b,c),$
where $F(a,b,c)$ is the following function
$\arctan\left(\frac{b-a-2}{\sqrt{d}}\right)-\arctan\left(\frac{b-a+2-2c}{\sqrt{d}}\right)+\arctan\left(\frac{-b+a-2}{\sqrt{d}}\right)-\arctan\left(\frac{-b+a+2-2c}{\sqrt{d}}\right).$
Notice that $F(a,b,c)=G(c,e)$ only depends on $c$ and e=bae=b-a because $d=4-4c-e^{2}$. It can be checked that $\partial_{c}G=\partial_{e}G=0$ so that $G$, and consequently $F$, is constant. Evaluating at $e=0$ we obtain that $F\equiv-\pi$.
To compute $\gamma_{-}$ it suffices to apply to $X$ the symmetry $(x,y)\mapsto(x,-y)$ that transforms the invariants $(a,b,c)$ into (a,b,c)(-a,-b,c) and the value $\gamma_{0}$ into $\gamma_{0}$. ∎
Figure 4: Local phase portrait of the vector field $(x^{2}+y^{2}+xy)\partial_{x}-(x+y)y\partial_{y}$ having first integral $\ln\!\left(y^{2}(2x^{2}+2xy+y^{2})\right)-2\arctan\!\left(\frac{x+y}{x}\right)$.
Consider the quadratic homogeneous vector field $X=(x^{2}+y^{2}+axy)\partial_{x}+(cx+by)y\partial_{y}$ with $d=4(1-c)-(a-b)^{2}>0$. Taking $\alpha=-1$ and $\omega=1$ we have that $\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0)}{f_{1}(x,0)}\frac{dx}{x}=\mathrm{PV}\int_{-1}^{1}\frac{cx}{x^{2}}dx=0$. Taking also $a=1$, b=1b=-1, c=1c=-1 (so that $d=4$) the transition map $\Pi_{-1}^{+1}$ is contractive for $y>0$ and expansive for $y<0$ in accordance with $\Pi_{-1}^{+1}(y)=e^{\gamma_{\pm}}y+\mathcal{F}_{1+\epsilon}$, where $\gamma_{\pm}=0\pm\pi\frac{(-2(-1)-(1-1)(-1))}{2}=\pm\pi$ depending on the sign $\pm$ of $y$, see Figure 4. $\square$
To finish this work we apply the previous two results to the study of the following family of degenerated singularities considered in [3]:
$Z_{\mu}:\left\{\begin{array}[]{l}\dot{x}=\beta x^{2}y+\alpha xy^{2}-\beta y^{3}-x^{4},\\ \dot{y}=4\beta xy^{2}+\alpha y^{3}+2x^{5},\end{array}\right.\qquad\mu=(\alpha,\beta)\in\mathbb{R}^{2}.$ (4)
According to [3, p. 189], the origin is monodromic for $Z_{\mu}$ if and only if $\beta>1/4$.
[…]
Figure 5: Blowup of the origin for the family $Z_{\mu}$ in the charts $\pi(x,u)=(x,ux)$ and $\pi(v,y)=(vy,y)$.
Blowing up the origin, see Figure 5, we have that
$Y_{\mu}=\frac{(x,ux)^{*}Z_{\mu}}{x^{2}}=(3\beta u^{2}+\beta u^{4}+ux+2x^{2})\partial_{u}+(\beta u+\alpha u^{2}-\beta u^{3}-x)x\partial_{x}$
and, assuming that $\beta>0$,
$X_{\mu}=\Big(\frac{x}{3\beta},\frac{y}{\sqrt{6\beta}}\Big)^{*}Y_{\mu}=\left(x^{2}+y^{2}+\frac{xy}{\sqrt{6\beta}}+\frac{x^{4}}{27\beta^{2}}\right)\partial_{x}+\left(\frac{x}{3}-\frac{y}{\sqrt{6\beta}}+\frac{\alpha x^{2}}{9\beta^{2}}-\frac{x^{3}}{27\beta^{2}}\right)y\partial_{y}$
is of the form (1) with $f_{1}(x,y;\mu)=1+\frac{x^{2}}{27\beta^{2}}$, $f_{2}(x,y;\mu)\equiv 1$, $a(\mu)=\frac{1}{\sqrt{6\beta}}$, $g_{1}(x,y;\mu)=\frac{1}{3}+\frac{\alpha x}{9\beta^{2}}-\frac{x^{2}}{27\beta^{2}}$, $g_{2}(y;\mu)\equiv b(\mu)=-\frac{1}{\sqrt{6\beta}}$ and $c(\mu)=\frac{1}{3}$.
Notice that $f_{1}(x,0;\mu)\neq 0$ for all $x\in\mathbb{R}$ and $d(\mu)=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}=\frac{2}{3}(4-\frac{1}{\beta})$ is positive if and only if $\beta>\frac{1}{4}$. By applying Theorem 2 we also deduce that the origin is monodromic if and only if $\beta>\frac{1}{4}$. In that case we can apply Theorem 4 and consider the limit (as $\eta\to 0^{+}$) transition map $\Pi_{-\infty}^{+\infty}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{\infty}(\{\beta>\frac{1}{4}\})$ of $X_{\mu}$, with
$\gamma_{\pm}(\mu)=\mathrm{PV}\int_{-\infty}^{+\infty}\frac{\frac{1}{3}+\frac{\alpha x}{9\beta^{2}}-\frac{x^{2}}{27\beta^{2}}}{1+\frac{x^{2}}{27\beta^{2}}}\frac{dx}{x}\mp\frac{\pi}{\sqrt{4\beta-1}}=\pi\left(\frac{\alpha}{\beta\sqrt{3}}\mp\frac{1}{\sqrt{4\beta-1}}\right).$
Assuming that $\beta>\frac{1}{4}$, the Poincaré return map $R(x;\mu)$ of $Z_{\mu}$ around $(0,0)$ is the composition of two transition maps $\Pi_{-\infty}^{+\infty}(x;\mu)=e^{\gamma_{\pm}(\mu)}x+\mathcal{F}_{1+\epsilon}^{\infty}(\{\beta>\frac{1}{4}\})$ of $X_{\mu}$, one for $x>0$ and the other for $x<0$, see Figure 5, so that $R(x;\mu)=e^{\gamma(\mu)}x+\mathcal{F}_{1+\epsilon}^{\infty}(\{\beta>\frac{1}{4}\})$ with $\gamma(\mu)=\gamma_{+}(\mu)+\gamma_{-}(\mu)=\frac{2\pi\alpha}{\beta\sqrt{3}}$, in agreement with [3, p. 189].
Notice that for any $\beta>\frac{1}{4}$ the origin is a center of $Z_{(0,\beta)}$ because it is reversible via $(x,y,t)\mapsto(-x,y,-t)$. Moreover, thanks to the uniform properties of the flat remainder class $\mathcal{F}_{1+\epsilon}^{\infty}$ we can assert that the ciclicity of the origin in the family (4) is zero at any $\mu=(\alpha,\beta)\in\mathbb{R}^{2}$ with $\beta>\frac{1}{4}$. This last assertion can not be deduced from the results in [3] because they are not uniform with respect to parameters.